OpenStudy (anonymous):

What is the degree measure of an arc 4 pi ft. long in a circle of radius 10 ft.?

OpenStudy (anonymous):

l = r * angle angle = l / r = 4pi/10 = 2pi/5 radians

OpenStudy (anonymous):

= 2pi/5 * 180/pi = 72 degrees

OpenStudy (anonymous):

i think i got the equation right this time!!!!